OpenStudy (anonymous):

Starting from 1.5 miles away, a car drives towards a speed checkpoint and then passes it. The car travels at a constant rate of 53 miles per hour. The distance of the car from the checkpoint is given by . At what times is the car 0.1 miles from the checkpoint? Calculate your answer in seconds

OpenStudy (paxpolaris):

Wait fill the blank... The distance of the car from the checkpoint is given by _______ ???

OpenStudy (anonymous):

the blank is..... d= 1.5 - 53t

OpenStudy (paxpolaris):

 Set d= 0.1 & solve for t

OpenStudy (paxpolaris):

\[0.1=1.5-53t\]

OpenStudy (paxpolaris):

\[\implies -53t=0.1-1.5=-1.4\]\[\implies 53t=1.4\]\[\implies t={1.4\over53}={14\over530}hours\]

OpenStudy (paxpolaris):

60 min in 1 hr .... 60 sec in 1 min \[t={14 \over 530} \cancel{hours} \cdot{60 \min \over 1\cancel{hour}}\cdot{60 \sec \over 1 \min}={14\cdot 60\cdot 60 \over 530}sec\]

OpenStudy (paxpolaris):

http://lmgtfy.com/?q=(14%2F530)+hours+in+seconds \[\huge t=\ \ 95.09...\ seconds\]